OpenStudy (anonymous):

WILL MEDAL AND FAN PLZ HELP data will be below Which equation is shown on the graph? A.y = 3x B.y = x – 3 C.y = x + 3 D.y = 2x – 3

OpenStudy (anonymous):

the graph pls ?

OpenStudy (anonymous):

when x = 3 y = 0 when x = 0 y = -3 x2-x1 = y2-y1 3-0/0- (-3) = 3 / 3 the slope is 1 y = 1x + b b is the y when x = 0 y = x - 3 hope that was helpful

OpenStudy (anonymous):

thx it was @Ogziii